# Anatomy of singlet-doublet dark matter relic: annihilation, co-annihilation, co-scattering, and freeze-in

### v1

@@ 3.2 Thermal dark matter relic via co-annihilation and
$+\left.\langle\sigma_{1222}v\rangle\left(Y_{1}Y_{2}-Y_{2}^{2}\frac{Y_{1}^{eq}}{Y_{2}^{eq}}\right)-\langle\sigma_{1211}v\rangle\left(Y_{1}Y_{2}-Y_{1}^{2}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)-\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)\right],$
$\frac{dY_{2}}{dT}$ $=$ $\frac{1}{3\mathcal{H}}\frac{ds}{dT}\left[\langle\sigma_{2200}v\rangle(Y_{2}^{2}-{Y_{2}^{eq}}^{2})-\langle\sigma_{1122}v\rangle\left(Y_{1}^{2}-Y_{2}^{2}\frac{{Y_{1}^{eq}}^{2}}{{Y_{2}^{eq}}^{2}}\right)+\langle\sigma_{1200}v\rangle(Y_{1}Y_{2}-Y_{1}^{eq}Y_{2}^{eq})\right.$
@@
$-\left.\langle\sigma_{1222}v\rangle\left(Y_{1}Y_{2}-Y_{2}^{2}\frac{Y_{1}^{eq}}{Y_{2}^{eq}}\right)+\langle\sigma_{1211}v\rangle\left(Y_{1}Y_{2}-Y_{1}^{2}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)+\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)\right],$
@@
where $Y_{i}^{eq}$ are the equilibrium abundances, $\mathcal{H}$ is the Hubble parameter, $\langle\sigma_{\alpha\beta\gamma\delta}v\rangle$ are the thermally averaged cross-sections for processes involving annihilation of particles of sectors $\alpha\beta\rightarrow\gamma\delta$. $\Gamma_{2\rightarrow 1}$ is the conversion term, which includes both the interaction rate of the co-scattering process ($\langle\sigma_{1020}v\rangle$) and the decay ($\Gamma_{D}$).
In the large mixing angle regime, only annihilation and co-annihilations are the dominating processes that decide the relic. In this case, the decoupling of the particles in both sectors happens in the same epoch, and these two equations simplify to the one-sector BE as given in Eq.(8). On the other hand, in the small SD mixing regime, annihilation and co-annihilation processes are sub-dominant. As a result Eq. (LABEL:eq:Y1) and (LABEL:eq:Y2) are simplified to
$\frac{dY_{1}}{dT}=-\frac{\Gamma_{2\rightarrow 1}}{\mathcal{H}T}\left[Y_{2}-Y_{1}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right],$ (15)
$\frac{dY_{2}}{dT}=\frac{s}{\mathcal{H}T}\left[\left<\sigma_{2200}v\right>\left(Y^{2}_{2}-{Y^{eq}_{2}}^{2}\right)+\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)\right].$ (16)
We have shown the evolution of the abundances of both sectors w.r.t. temperature for two different mixing angles: $\sin\theta=1.19\times 10^{-1}$ and $\sin\theta=1.19\times 10^{-5}$ in Fig. 2. We fixed the DM mass to 146.48 GeV and took the mass splitting to be 10.5 GeV and 8.84 GeV, respectively, to get the correct relic abundances in both cases. For the large mixing angle, we see that both sectors remain in equilibrium for around $T\sim 6$ GeV in the left panel of Fig. 2. In this case, co-annihilation is the dominant process, which is deciding the relic. When we decrease the mixing angle by four orders of magnitudes, the co-annihilation and annihilation processes, which depend on the mixing angle as $\sin^{2}\theta$ and $\sin^{4}\theta$ respectively, become negligible. We see that the sector 1 particle goes out of equilibrium very early at around $T\sim 8$ GeV, whereas the sector 2 particles are still in equilibrium due to their gauge interactions with the SM bath in right panel of Fig. 2. The sector 1 particle goes out of equilibrium with a larger abundance, then upscatters with the SM particles to the sector 2 particles, which is the co-scattering process. Thus, the relic of sector 1 comes down to the correct ballpark.
@@
Figure 2: Evolutions of sector 1 and sector 2 particle abundances for $M_{\rm DM}$=146.48 (146.48) GeV, $M_{\psi}=156.98~{}(155.32)$ GeV, and $\sin\theta=1.19\times 10^{-1}~{}(1.19\times 10^{-5})$ are shown in left (right) plot which gives the DM relic of $\Omega_{\rm DM}h^{2}\simeq 0.12$.
We now explore the regime of the co-scattering effects in the SDDM scenario, which was not taken into account in earlier works [9, 8, 10, 11, 12, 13, 14, 15, 16, 17, 18]. To explore how the relic density of DM depends on our model parameters, we performed several analyses and examined the allowed parameter space.
@@
Figure 3: Variation of DM relic density w.r.t. DM mass with the fixed mass splitting of $51<\Delta{M}<$ 5 GeV for $10^{-2}<\sin\theta<10^{-1}$ [left] and $10^{-5}<\sin\theta<10^{-4}$ [right] considering co-scattering processes (red) and without considering co-scattering processes (blue).
We study the relic density of DM considering co-scattering and without co-scattering with different sets of our model parameters. In Fig. 3, we have kept the mass splitting in the range of 1 GeV to 5 GeV and then varied the mixing angle in two sets. In the left pannel of Fig. 3, the mixing angle is varied in between $10^{-2}-10^{-1}$. Due to smaller mass splitting and the large mixing angle, the co-annihilations are the most dominating processes that determine the relic of the DM. Co-scattering effects are sub-dominant in this region as the mixing angle is large enough to keep the DM in the thermal equilibrium for a longer period. The freeze-out via co-annihilation processes finally decides the relic of the DM in this case. We see that the relic density points obtained considering the co-scattering and without co-scattering coincide, validating the co-annihilation regime. Now, keeping the mass splitting in the same range, we decrease the mixing angle to a range of $10^{-5}-10^{-4}$. Now, because of this small mixing angle, the annihilation/co-annihilation processes become subdominant, and the DM gets chemically decoupled, leading to a larger DM abundance. These points are shown with the blue dots in the right panel of Fig. 3. Now, when we consider the co-scattering effects, the dynamics of the freeze-out of DM change drastically. The DM can now scatter with SM particles to a higher mass state, $\psi^{\pm},\chi_{0}$, and SM particles, known as co-scattering processes or conversion-driven processes, followed by annihilation of doublets to SM particles. These processes help the DM to bring down its relic density to the correct ballpark. These points are shown in red color in the right panel of Fig. 3. We see that the co-scattering effects are significant in the small mixing angle regime while they are sub-dominant for the large mixing angles.
@@
Figure 4: Variation of DM relic density w.r.t DM mass with the fixed mass splitting of $105<\Delta{M}<$ 10 GeV for $10^{-2}<\sin\theta<10^{-1}$ [left] and $10^{-5}<\sin\theta<10^{-4}$ [right] considering co-scattering processes (red) and without considering co-scattering processes (blue).
We now move to Fig. 4, where we have varied the mass splitting in the range from 5 to 10 GeV, while the SD mixing is kept in the same range as in Fig. 3. We observe the dependency of the DM relic on the SD mixing angles remains similar as in Fig. 3. When we increase the mass splitting, the co-annihilation or co-scattering cross becomes smaller, which leads to larger DM abundance. This is clearly visible from the left and right panel of Fig. 4.
@@
Figure 5: Correct relic parameter space in the plane of $\Delta{M}-M_{\rm DM}$ excluding co-scattering (left) and including co-scattering (right) with different ranges of mixing angles as mentioned in the plot.

### v2

@@ 3.1 Thermal production of dark matter
$+\left.\langle\sigma_{1222}v\rangle\left(Y_{1}Y_{2}-Y_{2}^{2}\frac{Y_{1}^{\rm eq}}{Y_{2}^{\rm eq}}\right)-\langle\sigma_{1211}v\rangle\left(Y_{1}Y_{2}-Y_{1}^{2}\frac{Y_{2}^{\rm eq}}{Y_{1}^{\rm eq}}\right)-\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{\rm eq}}{Y_{1}^{\rm eq}}\right)\right],$
$\frac{dY_{2}}{dT}$ $=$ $\frac{1}{3\mathcal{H}}\frac{ds}{dT}\left[\langle\sigma_{2200}v\rangle(Y_{2}^{2}-{Y_{2}^{\rm eq}}^{2})-\langle\sigma_{1122}v\rangle\left(Y_{1}^{2}-Y_{2}^{2}\frac{{Y_{1}^{\rm eq}}^{2}}{{Y_{2}^{\rm eq}}^{2}}\right)+\langle\sigma_{1200}v\rangle(Y_{1}Y_{2}-Y_{1}^{\rm eq}Y_{2}^{\rm eq})\right$.
@@
$-\left.\langle\sigma_{1222}v\rangle\left(Y_{1}Y_{2}-Y_{2}^{2}\frac{Y_{1}^{\rm eq}}{Y_{2}^{\rm eq}}\right)+\langle\sigma_{1211}v\rangle\left(Y_{1}Y_{2}-Y_{1}^{2}\frac{Y_{2}^{\rm eq}}{Y_{1}^{\rm eq}}\right)+\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{\rm eq}}{Y_{1}^{\rm eq}}\right)\right],$
@@
where $Y_{i}^{\rm eq}\left(=\frac{n_{i}^{\rm eq}}{s}\right)$ are the equilibrium abundances, $\mathcal{H}$ is the Hubble parameter, the entropy density, $s=\frac{2\pi^{2}}{45}g_{*s}T^{3}$, $\langle\sigma_{\alpha\beta\gamma\delta}v\rangle$ are the thermally averaged cross-sections for processes involving annihilation of particles of sectors $\alpha\beta\rightarrow\gamma\delta$, which is given by [52]
$\langle\sigma_{\alpha\beta\gamma\delta}v\rangle=\frac{T}{8m_{\alpha}^{2}m_{\beta}^{2}K_{2}(\frac{m_{\alpha}}{T})K_{2}(\frac{m_{\beta}}{T})}\int_{(m_{\alpha}+m_{\beta})^{2}}^{\infty}\sigma_{\alpha\beta\rightarrow\gamma\delta}(s)\big{(}s-(m_{\alpha}+m_{\beta})^{2}\big{)}\sqrt{s}K_{1}\bigg{(}\frac{\sqrt{s}}{T}\bigg{)}ds,$
and $\Gamma_{2\rightarrow 1}$ is the conversion term, which includes both the interaction rate of the co-scattering process and the decay defined as
$\Gamma_{2\rightarrow 1}=\Gamma_{\Psi\rightarrow\chi,\rm SM}\frac{K_{1}(M_{\Psi}/T)}{K_{2}(M_{\Psi}/T)}+\langle\sigma_{2010}v\rangle n^{\rm eq}_{\rm SM},$ (3.8)
where $\Gamma_{\Psi\rightarrow\chi,\rm SM}$ includes two body decays like $\Gamma_{\chi_{0}\rightarrow\chi_{1}H},\Gamma_{\chi_{0}\rightarrow\chi_{1}Z},\Gamma_{\psi^{\pm}\rightarrow\chi_{1}W^{\pm}}$, and three body decay of the doublets as $\Gamma_{\psi^{\pm}\rightarrow\chi_{1}l^{\pm}\nu_{l}}$ [^3] The decay mode $\Gamma_{\psi^{\pm}\rightarrow\chi_{1}\pi^{\pm}}$ is very suppressed. This is because of $\Delta M\geq 1$ GeV.. The decay rates are given in Appendix C. $\langle\sigma_{2010}v\rangle$ is the thermally averaged cross-section of the co-scattering processes with $n_{\rm SM}^{\rm eq}=0.238\times s$ [53].
3.1.1 Thermal DM relic in the large mixing regime via annihilation and co-annihilation
We are now well equipped to calculate the thermal relic in the both small and large $\sin\theta$ regime. As discussed before for $\sin\theta\sim\mathcal{O}(10^{-2})$, the dark sector particles decouple chemically at the same epoch. To validate this assumption of chemical equilibrium among the dark sector particles we define a new parameter $\rho_{1}$ as
$\rho_{1}=\frac{\Omega_{1s}h^{2}}{\Omega_{2s}h^{2}(\rm no\leavevmode\nobreak\ coscattering)},$ (3.9)
where $\Omega_{1s}h^{2}$ is the relic density obtained by solving Eq. (3.1), and $\Omega_{2s}h^{2}$(no co-scattering) is calculated by solving Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2) excluding co-scattering processes [^4] We use micrOMEGAs 6.1.15[53, 42] to solve the Boltzmann equations, where darkOmega function is used to estimate $\Omega_{1s}h^{2}$ and darkOmegaN is used to estimate $\Omega_{2s}h^{2}$(no co-scattering) while excluding the co-scattering processes. This can be achieved by using the flag ExcludedForNDM="2010", while using darkOmegaN function.. In particular we have switched off $\langle\sigma_{2010}v\rangle n^{\rm eq}_{\rm SM}$ in Eq. (3.8). In Eq. (3.9), $\rho_{1}=1$ implies that all the dark sector particles decouple at the same epoch. On the other hand, $\rho_{1}<1$ implies that the dark sector particles are decoupling at different epochs. When $\sin\theta$ is small, the singlet decouples earlier than the doublet and produces a relatively larger relic in absence of the co-scattering processes. This implies that $\Omega_{2s}h^{2}({\rm no\leavevmode\nobreak\ coscatttering})>\Omega_{1s}h^{2}$ Therefore, one expects $\rho_{1}\leq 1$.
Figure 1: Allowed parameter space of DM for which DM is in equilibrium with SM bath and the $\mathcal{Z}_{2}$-odd doublet. All the points in the plots (left and right) satisfy the required relic density. The grey-shaded region in both the plots shows the LEP bound [54] on the heavy fermion doublet.
We have presented the parameter space in Fig. 1. In this plot, all the points satisfy the correct relic density of DM, and we have used $0.99<\rho_{1}\leq 1$ to assess the validity and domain of applicability of the Eq. (3.1). In Fig. 1 (left), we have shown all the points in the plane of $\sin\theta$ and $M_{\rm DM}$, and in the right plot, we projected the same points in the plane of $\Delta{M}$ and $M_{\rm DM}$. DM mass is varied in the range of 110001-1000 GeV, $\Delta{M}$ is varied in the range of 110001-1000 GeV, and $\sin\theta$ is varied in the range of ($0.5-10^{-4}$). From Fig. 1 (left), we observe that the value of $\sin\theta$ can be as small as $\sim 5\times 10^{-4}$ for $M_{\rm DM}\sim 250$ GeV. As shown in Fig. 1 (right), we exactly reproduced all the points that satisfy correct relic density in the plane of $\Delta{M}$ vs $M_{\rm DM}$ as given in [6]. The pattern shown in the right plot arise from the interplay between annihilation and co-annihilation processes. In the region with small mass splitting ($\Delta{\rm M}$ $\lesssim 15$ GeV), the relic density satisfying points can be attributed to dominant co-annihilation channels. As the dark matter mass increases, the thermally averaged cross section decreases; consequently, a smaller mass splitting is required to enhance the co-annihilation rate and achieve the correct relic abundance. The blank region below the colored band for $M_{\rm DM}<700$ GeV corresponds to an under-abundant relic density, primarily due to enhanced co-annihilation cross sections at lower masses. Around $\Delta M\sim$ 50 GeV, the contributions from co-annihilation and annihilation processes become comparable. Above this threshold, annihilation channels dominantly determine the relic density. In this regime, the cross section of Higgs-mediated annihilation processes grows with the Yukawa coupling, $y$, where $y\propto\Delta{M}\sin\theta$. Therefore, with increasing dark matter mass for a fixed $\sin\theta$, a corresponding increase in $\Delta{M}$ is necessary to provide the observed relic abundance. Consequently, the blank region on the right side of the plot represents parameter space yielding an over-abundant relic density while the region to the left is in the correct ball park.
Thus we tested the validity of Eq. (3.1) for $\sin\theta\in[10^{-4},0.5]$. In other words, solving Eq. (3.1) is equivalent to solving Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2) keeping $\langle\sigma v\rangle_{2010}$ switched off. This confirms the earlier assumption [6] that singlet and doublet decouple chemically at the same epoch in the large $\sin\theta$ limit. In this region of parameter space even if we do not switch off the co-scattering term in Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), the results remain unchanged. See more discussion below.
3.1.2 Thermal dark matter relic via the conversion-driven processes in the small $\sin\theta$ regime
We have already demonstrated in Fig. 1 (left) the parameter space for the correct relic of DM in $\sin\theta$-$M_{\rm DM}$ plane, where the singlet and the doublet freeze out at the same epoch through annihilation and co-annihilation processes. Beyond this regime, the singlet may reach equilibrium. However, in this case, the singlet and doublet will decouple at different epochs. Due to small mixing, the singlet will decouple early, while the doublet components will remain in equilibrium for a longer period due to their gauge interactions. Although the singlet decouples early with a larger abundance, the final relic of the DM can be brought to the correct ballpark in the presence of additional conversion-driven processes [40, 41].
As discussed in section 3.1.1, in the large SD mixing angle limit ($\mathcal{O}(10^{-4})<\sin\theta<0.5$) only annihilation and co-annihilation are the dominating processes that decide the DM relic. Had we been added the co-scattering processes to Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), the result would not have changed. This can be understood by analyzing Eq. (3.8), where the second term contributes to the effective co-scattering processes. At this point we note that in the large $\sin\theta$ limit, the impact of the co-scattering processes are similar in magnitude to that of the decay and inverse decay processes given in Eq. (3.8). The large co-scattering processes can interchange the sector 1 particles to sector 2 particles efficiently. If the sector 2 particles are already in equilibrium, then they can annihilate to SM particles and deplete the dark matter number density through annihilation and co-annihilation processes $\psi^{+}\psi^{-}\rightarrow\rm\rm SM\leavevmode\nobreak\ SM$ which is independent of the SD mixing angle. On the other hand, if the singlet and doublet decoupled already (the processes 1100,1200 and 2200 decoupled) then the large conversion-driven processes can not alter the total relic, $Y_{1}+Y_{2}$, even if these processes remain in equilibrium for a longer period. As a result, the impact of the co-scattering processes is not felt in the large mixing angle limit. In this case, the decoupling of particles in both sectors happens in the same epoch and the Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2) simplify to the one-sector BE as given in Eq. (3.1).
On the other hand, in the small SD mixing regime, annihilation and co-annihilation processes are sub-dominant. In this limit, Eq. (LABEL:eq:Y1) and (LABEL:eq:Y2) are simplified to
$\frac{dY_{1}}{dT}=-\frac{\Gamma_{2\rightarrow 1}}{\mathcal{H}T}\left[Y_{2}-Y_{1}\frac{Y_{2}^{\rm eq}}{Y_{1}^{\rm eq}}\right],$ (3.10)
$\frac{dY_{2}}{dT}=\frac{s}{\mathcal{H}T}\left[\left<\sigma_{2200}v\right>\left(Y^{2}_{2}-{Y^{\rm eq}_{2}}^{2}\right)+\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{\rm eq}}{Y_{1}^{\rm eq}}\right)\right].$ (3.11)
@@ 3.1.2 Thermal dark matter relic via the conversion-driven processes in the small $\sin\theta$ regime
Figure 2: Left: Evolutions of sector 1 and sector 2 particle abundances for $M_{\rm DM}$=146.48 GeV, $\Delta M=10.5$ GeV, and $\sin\theta=1.19\times 10^{-1}$ are shown for correct DM relic density $\Omega_{\rm DM}h^{2}\simeq 0.12$. Right: The comparison of the interaction rates w.r.t Hubble is shown as a function of temperature for the parameters given in the left panel.
Figure 3: Left: Evolutions of sector 1 and sector 2 particle abundances for $M_{\rm DM}$=146.48 GeV, $\Delta{M}=8.4$ GeV, and $\sin\theta=1.19\times 10^{-5}$ are shown for correct DM relic density $\Omega_{\rm DM}h^{2}\simeq 0.12$. Right: The comparison of the interaction rates w.r.t Hubble is shown as a function of temperature for the parameters given in the left panel.
We now explore the regime of the conversion-driven processes, including co-scatterings, decays and inverse decays, whose impact on the SDDM relic was not studied in earlier works [9, 6, 10, 11, 12, 13, 14, 15, 16, 17, 18]. We consider two benchmark points BP1, BP2: one in the large mixing limit (BP1 in Fig. 2) and another in small mixing limit (BP2 in Fig. 3).
BP1 ($\sin\theta=0.119$): In the left panel of Fig. 2, we depict the evolution of the abundances of both dark sector species as a function of temperature for mixing angles: $\sin\theta=1.19\times 10^{-1}$. The dark matter mass is fixed at $146.48\leavevmode\nobreak\ \mathrm{GeV}$, and $\Delta{M}=10.5\leavevmode\nobreak\ \mathrm{GeV}$, selected to obtain the correct relic abundance. The comparison among relevant interaction rates is displayed in the right panel of Fig. 2. Here, we observe that the annihilation rate of the singlet component (1100 processes) falls below the Hubble rate first, followed by the decoupling of co-annihilation processes (1200 processes). Both sectors remain in chemical equilibrium down to $T\sim 6\leavevmode\nobreak\ \mathrm{GeV}$, as seen in the left panel of Fig. 2. On the other hand, the rate of conversion-driven processes (decay, inverse decay and co-scattering) remain significantly larger than the Hubble rate during this period. These processes can efficiently deplete the number density of dark matter by efficiently converting the sector 1 particles to sector 2 particles followed by annihilation to SM particles (2200 processes). Once the 2200 processes decouple (shown by the red line in right panel of Fig. 2) around $T\sim 5$ GeV, the number density is settled to its final value as seen in the left panel of Fig. 2.
BP2 ($\sin\theta=1.19\times 10^{-5}$): In the left panel of Fig. 3, we depict the evolution of the abundances of both dark sector species as a function of temperature for mixing angles: $\sin\theta=1.19\times 10^{-5}$. The DM mass is taken to be 146.48 GeV, and $\Delta{M}=8.84\leavevmode\nobreak\ \mathrm{GeV}$ to get the correct relic density. Since the rate of sector 1 annihilation processes (1100) is proportional to $\sin^{4}\theta$, these processes are negligible in comparison to co-annihilation processes (1200) for this BP as the rate of co-annihilation processes are proportional to $\sin^{2}\theta$. As shown in the right panel of Fig. 3, the co-annihilation processes decouple at approximately $T\sim 20\leavevmode\nobreak\ \mathrm{GeV}$. As a result the chemical decoupling of sector 1 particles happens with a relatively larger abundance, while sector 2 particles remain in equilibrium due to their gauge interactions, as shown in the left panel of Fig. 3. Note that in the small mixing angle limit, the decay and inverse decay processes are subdominant in comparison to co-scattering processes. This is evident from the right panel of Fig. 3. It is worth mentioning that, in this case, processes involving light SM fermions contribute to the co-scattering processes significantly, e.g. $\chi_{1}e^{-}\rightarrow\psi^{-}\nu_{e},\chi_{1}\nu_{e}\rightarrow\psi^{+}e^{-}$ etc. In the presence of efficient co-scattering, sector 1 particle can upscatter with SM particles into sector 2, followed by annihilation to SM via 2200 processes. Once the 2200 processes decouple, the relic settles to its final value even if the rate of co-scattering processes continue to be in equilibrium for a longer period. For this BP the 2200 processes decouple at $T\sim 5$ GeV (see red line in the right panel of Fig. 3.). As a result the relic settles around $T\sim 5$ GeV as seen from the left panel of Fig. 3.
To explore in details the impact of co-scattering on the relic density of DM, we performed several analyses and examined the allowed parameter space by solving Eq. (LABEL:eq:Y1) and (LABEL:eq:Y2).
@@
Figure 4: Variation of DM relic density w.r.t. DM mass with the fixed mass splitting of $1\leavevmode\nobreak\ \rm GeV\leq\Delta{M}\leq 5$ GeV for $10^{-2}\leq\sin\theta\leq 10^{-1}$ [left] and $10^{-5}\leq\sin\theta\leq 10^{-4}$ [right]. For the blue colored points, we have solved Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), by switching off the co-scattering processes, while for the red points we have considered all the processes given in Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2).
In Fig. 4, we have kept the SD mass splitting in the range of 1 GeV to 5 GeV and then varied $\sin\theta$ in two sets. In the left panel of Fig. 4, $\sin\theta$ is varied in the range $10^{-2}-10^{-1}$, while in the right panel it is varied in the range $10^{-5}-10^{-4}$ . In the left panel, due to smaller mass splitting and the large mixing angle, the co-annihilations are the most dominant processes that determine the relic of the DM. We see that the red points (obtained by solving Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2)) and blue points (obtained by solving Eq (LABEL:eq:Y1) and (LABEL:eq:Y2), but keeping the co-scattering processes off) coincide with each other [^5] In Fig. 4, and Fig. 5 without co-scattering points are obtained using the function darkOmegaN with the flag ExcludedForNDM="2010", whereas with co-scattering points are obtained using darkOmegaN function without any flag.. This implies that in the large $\sin\theta$ limit, the impact of co-scattering is not felt. This is because, for large $\sin\theta$, the sector 1 (singlet) and sector 2 (doublet) decouple at the same epoch. In other words, the processes 1100, 1200 and 2200 decouple close to each other. In this case, even if the rate of co-scattering processes are large, they are not affecting the overall relic abundance ($Y_{1}+Y_{2}$) as the sector 1 and sector 2 particles are already decoupled from the thermal bath. In this case, the conversion-driven processes can at most interchange the number density of the sector 1 and sector 2 particles. However, the total dark sector relic, $Y_{1}+Y_{2}$ will be remain unchanged.
In the right panel of Fig. 4, because of the small mixing angle ($10^{-5}\leq\sin\theta\leq 10^{-4}$), the annihilation/co-annihilation processes become subdominant. In the absence of co-scattering processes, the DM gets chemically decoupled with a larger DM abundance as shown by the blue points. When we include the co-scattering effects, the dynamics of the freeze-out of DM change drastically. In particular these processes further deplete the DM relic density by efficiently converting the sector 1 particles to sector 2 particles followed by annihilation to SM bath until the 2200 processes are decoupled. This is shown by the red color points. We see that the co-scattering effects are significant in the small mixing angle regime while their effects are not prominent for the large mixing angles even if the rate of these processes are large.
@@
Figure 5: Variation of DM relic density w.r.t DM mass with the fixed mass splitting of $5\leavevmode\nobreak\ {\rm GeV}\leq\Delta{M}\leq 10$ GeV for $10^{-2}\leq\sin\theta\leq 10^{-1}$ [left] and $10^{-5}\leq\sin\theta\leq 10^{-4}$ [right]. For the blue colored points, we have solved Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), by switching off the co-scattering processes, while for the red points we have considered all the processes given in Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2).
We now move to Fig. 5, where we have varied the mass splitting in the range from 5 GeV to 10 GeV, while the SD mixing is kept in the same range as in Fig. 4. We observe that the dependency of the DM relic on the SD mixing angles remains similar as in Fig. 4. When we increase the mass splitting, the co-annihilation and co-scattering crosssection becomes smaller, which leads to larger DM abundance in comparison to Fig. 4. This is clearly visible from the left panel of Fig. 5. Now we compare the red colored points in the right panels of Fig. 4 and 5 where the $\Delta{M}$ increases in Fig. 5 compared to Fig. 4 while all other parameters remain same. A larger $\Delta M$ reduces the co-annihilation and co-scattering effects. Moreover, in the small $\sin\theta$ limit co-scattering effects always dominate over decay inverse decay. As a result we get an overall increment in the dark matter relic abundance. This can be easily seen by looking at the red colored points in the right panel of Fig. 4 and 5. We then compare the blue colored points (co-scattering effects are not included while decay and inverse decay terms are present, see Eq. (3.8)) in the right panel of Fig. 4 and 5. As the $\Delta{M}$ increases in Fig. 5 w.r.t Fig. 4, the co-annihilation effects decrease. On the other hand, the decay and inverse decay effects increase as they are proportional to $(\Delta{M})^{5}$ (see Eq. (C.6)). The latter processes help in depleting the relic for a longer period. As a result we get an overall reduced relic even though there is a small relic enhancement due to the reduced co-annihilation effect.
@@
Figure 6: Right: The correct relic density is shown in the plane $\Delta{M}-M_{\rm DM}$ with different ranges of mixing angles as provided in the figure inset by solving Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2). Left: Same as in the right panel, except that the co-scattering effects are switched off.
In Fig. 6, we present the points that yield the correct DM relic density in the $\Delta M$-$M_{\rm DM}$ plane by solving Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2). In the left plot, we show the points that give the correct relic density when co-scattering processes are switched off. In this case, the relic density calculation considers only annihilation, co-annihilation, decay and inverse decay processes. The color coding represents different values of $\sin\theta$, ranging from 0.5 to $5\times 10^{-4}$, as indicated in the plot. For $\sin\theta\geq 0.05$, the singlet and doublet components decouple almost at the same epoch. Therefore, our results, obtained by solving Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), are consistent with previous findings [6], obtained by solving Eq (3.1) and depicted in the right panel of Fig. 1. For $\sin\theta<0.05$, singlet and doublet components decouple at different epochs. As a result we find new parameter space shown in magenta and cyan colored points in comparison to the right panel of Fig. 1.